OpenStudy (anonymous):

37. Write y = –2/5x – 3 in standard form using integers. (1 point) 2x + 5y = –15 2x + y = –3 2x + 5y = –3 x + y = –15

OpenStudy (anonymous):

Standard form is: ax + by = c So, y + 3 = (-2/5)x 5y + 15 = -2x 2x + 5y = -15

OpenStudy (anonymous):

You're moving right along!